1 solid is a sphere and has a radius of 6 in the other solid is a cylinder with a radius of 6 in and a height of 6 in what is the difference the volumes have
in

Found 3 solutions by josgarithmetic, math_tutor2020, ikleyn:
Answer by josgarithmetic(39617)

You know or can find the formulas for volume of sphere and cylinder?

Answer by math_tutor2020(3816)


V = (4/3)*pi*r^3 .... volume of a sphere
V = (4/3)*pi*6^3
V = 288pi
The sphere has an exact volume of 288pi cubic inches.


V = pi*r^2*h .... volume of a cylinder
V = pi*6^2*6
V = 216pi
The cylinder has an exact volume of 216pi cubic inches

The difference of the volumes is
sphere - cylinder = 288pi - 216pi = 72pi cubic inches

Extra info:
The ratio of the volumes is
cylinder/circle = (216pi)/(288pi) = 3/4
meaning that the cylinder is exactly 3/4 the volume of the sphere
cylinder volume = (3/4)*(sphere volume)
This only works if r = h, and r is the same for both the sphere and cylinder.


Answer by ikleyn(52777)

.


This discovery is attributed to Archimedes, a great ancient Greece mathematician, physicist and engineer (c. 287 - c. 212 BC).


He was first who discovered that in this situation

        volume of the cylinder is equal to the volume of the inscribed sphere PLUS the volume of the inscribed cone with the same base.


The legend says that Archimedes bequeathed to carve this formula on his grave.


About Archimedes, read this Wikipedia articles

https://en.wikipedia.org/wiki/Archimedes

https://simple.wikipedia.org/wiki/Archimedes